# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ V Conclusion
VI Acknowledgements
A.J.D. acknowledges support from the NIH/CU Molecular Biophysics Program and the NIH Biophysics Training Grant T32 GM145437. N.L.C. acknowledges the Arnold and Mabel Beckman Foundation Beckman Scholars Program (BSP2022). W.C.P. acknowledges support from Cottrell Scholar Award CS-CSA-2023-077 sponsored by the Research Corporation for Science Advancement and from Integrated Research-Education Grant KA2022-129526 from the Charles E. Kaufman Foundation of the Pittsburgh Foundation. R.J.R. acknowledges financial support from grant CS-CSA-2024-098 from Research Corporation for Science Advancement. A.M.C. acknowledges the start-up funds from the University of Colorado Boulder. The NMR spectrometer was funded by support from the National Science Foundation (CHE-0957839). The authors thank Dr. Paolo Suating for help in collecting experimental NMR spectra and Prof. Meredith Borden for feedback on the article.

### v2

@@ V Conclusion
VI Associated Content
Our materials are freely available at the following GitHub repository: https://github.com/Montoya-Castillo-Group/FourierTransforms. Students can run each .ipynb file by manually uploading them to Google Colab.
VII Acknowledgments
A.J.D. acknowledges support from the NIH/CU Molecular Biophysics Program and the NIH Biophysics Training Grant T32 GM145437. N.L.C. acknowledges the Arnold and Mabel Beckman Foundation Beckman Scholars Program (BSP2022). W.C.P. acknowledges support from Cottrell Scholar Award CS-CSA-2023-077 sponsored by the Research Corporation for Science Advancement and from Integrated Research-Education Grant KA2022-129526 from the Charles E. Kaufman Foundation of the Pittsburgh Foundation. R.J.R. acknowledges financial support from grant CS-CSA-2024-098 from Research Corporation for Science Advancement. A.M.C. acknowledges the start-up funds from the University of Colorado Boulder. The NMR spectrometer was funded by support from the National Science Foundation (CHE-0957839). The authors thank Dr. Paolo Suating for help in collecting experimental NMR spectra and Profs. Meredith Borden, Matthew Pons, and David Schmitz for feedback on the article. We additionally thank Prof. David Horner for insightful discussions and feedback on the final manuscript.
VIII Appendix
VIII.1 Rewriting the Fourier sine and cosine series in terms of complex exponentials
We start with the Fourier ‘sine and cosine’ series of a periodic, smooth function $f$, given by
$f(t)=a_{0}+\sum_{n=1}^{\infty}a_{n}\cos\left(\omega_{n}t\right)+\sum_{n=1}^{\infty}b_{n}\sin\left(\omega_{n}t\right),$ (8)
where
$a_{n}=\begin{cases}\dfrac{1}{2T}\displaystyle\int_{-T}^{T}f(t)\,{\rm d}t,&n=0;\\[15.00002pt] \dfrac{1}{T}\displaystyle\int_{-T}^{T}f(t)\cos(\omega_{n}t)\,{\rm d}t,&n\geq 1,\end{cases}$ (9)
and
$b_{n}=\frac{1}{T}\int_{-T}^{T}f(t)\sin(\omega_{n}t)\,{\rm d}t,$ (10)
for $n\geq 1$. Then, we insert the identities
$\cos(\theta)=\frac{e^{i\theta}+e^{-i\theta}}{2}\,\,\,\text{ and }\,\,\,\sin(\theta)=\frac{e^{i\theta}-e^{-i\theta}}{2i}$
into Eq. (8) to obtain
$\begin{split}f(t)&=\sum_{n=0}^{\infty}\left[a_{n}\cos(\omega_{n}t)+b_{n}\sin(\omega_{n}t)\right]\\ &=\!\sum_{n=0}^{\infty}\!\bigg{[}\dfrac{a_{n}}{2}\big{(}e^{i\omega_{n}t}\!+e^{-i\omega_{n}t}\big{)}+\dfrac{b_{n}}{2i}\big{(}e^{i\omega_{n}t}-e^{-i\omega_{n}t}\big{)}\bigg{]}\\[1.00006pt] &=\!\sum_{n=0}^{\infty}\bigg{[}\dfrac{1}{2}\left(a_{n}-ib_{n}\right)e^{i\omega_{n}t}+\dfrac{1}{2}\left(a_{n}+ib_{n}\right)e^{-i\omega_{n}t}\bigg{]}\\[1.00006pt] &=\!\sum_{n=0}^{\infty}\dfrac{1}{2}\left(a_{n}-ib_{n}\right)e^{i\omega_{n}t}\!+\!\!\sum_{n=-1}^{-\infty}\dfrac{1}{2}\left(a_{-n}+ib_{-n}\right)e^{i\omega_{n}t}\\[1.00006pt] &\equiv\frac{1}{T}\sum_{n=-\infty}^{\infty}\hat{f}_{n}e^{i\omega_{n}t},\end{split}$
where we have used the fact that $\omega_{n}=2\pi n/T$ re-indexes the sum (i.e., $n\rightarrow-n$) in the penultimate equality, allowing one to define the Fourier coefficients $\{\hat{f}_{n}\}$ as
$\hat{f}_{n}=\begin{cases}\dfrac{T}{2}(a_{n}-ib_{n}),&\,\,\,n>0;\\[10.00002pt] Ta_{0},&\,\,\,n=0;\\[10.00002pt] \dfrac{T}{2}(a_{n}+ib_{n}),&\,\,\,n<0.\end{cases}$ (11)
Inserting Eqs (9) and Eq. (10) into Eq. (11), one can write the general expression
$\hat{f}_{n}=\int_{-T}^{T}f(t)e^{-i\omega_{n}t}\,{\rm d}t.$ (12)
VIII.2 Fourier series as period length approach infinity
Here, we derive Eqs. (4) and (5) from Eqs. (2) and (3). The key is to think of the function $f$ as being periodic with period length approaching infinity, i.e., $T\rightarrow\infty$. Consequently, as $T\rightarrow\infty$ the difference between $\omega_{n}$ and $\omega_{n+1}$ becomes infinitesimally small. Thus, one can invoke the limit definition of the integral to connect the summation over the discrete index $n$ in the Fourier series to an integral over a continuous variable in the FT.
To start, we introduce a new variable to keep track of the difference between consecutive frequencies, $\Delta=\omega_{n+1}-\omega_{n}=2\pi/T$. Then, we insert this into Eq. (3) and obtain
$f(t)=\frac{1}{2\pi}\!\sum_{n=-\infty}^{\infty}\!\frac{2\pi}{T}\hat{f}_{n}\,e^{i\omega_{n}t}=\!\frac{1}{2\pi}\!\sum_{n=-\infty}^{\infty}\!\!\Delta\,\hat{f}_{n}\,e^{i\omega_{n}t}.$
As the period length of $f$ becomes infinite, $\Delta=\omega_{n+1}-\omega_{n}\rightarrow 0$ and thus $\omega_{n}$ becomes a continuous variable, $\omega$. It also follows that $\hat{f}_{n}\rightarrow\hat{f}(\omega)$, $\Delta\rightarrow{\rm d}\omega$, and $\sum_{n}\Delta\rightarrow\int{\rm d}\omega$. Mathematically,
$f(t)\!\equiv\frac{1}{2\pi}\int_{-\infty}^{\infty}\hat{f}(\omega)e^{i\omega t}\,{\rm d}\omega\!=\!\lim_{\Delta\rightarrow 0}\sum_{n=-\infty}^{\infty}\hat{f}_{n}e^{i\omega_{n}t}\Delta\!\,.$ (13)
Similarly, since $\omega_{n}$ becomes continuous as $T\rightarrow\infty$, we see that
$\hat{f}(\omega)\equiv\lim_{T\rightarrow\infty}\hat{f}_{n}=\lim_{T\rightarrow\infty}\int_{-T}^{T}f(t)e^{-i\omega_{n}t}\,{\rm d}t.$ (14)